In 2017 Kelly calculated a ($18,821) NOL on her Form 1045 Schedule A. She wants to carryback her 2017 NOL to 2015. In 2015 Kelly filed single, had no dependents, took the standard deduction and her AGI reported on her Form 1040, line 37, was $46,467. Kelly’s 2015 income tax was $7,065. Using Form 1045, pages 1 and 2, carryback Kelly’s 2017 NOL to 2015 and calculate her refund
Answer #1
AGI of 2015- $ 46467
(carryback )NOL of 2017 - (18821)$
Net - $ 27646
Tax bracket :
$9,226 to $37,450 $922.50 plus 15% of the amount over $9,225
922.5+(27646-9225)*15/100= 922.5+2763 = $ 3685
Already paid for 2015 = $7065
Refund =7065-3685 $                                          3,380.00